# Making FETCH! Happen: Finding Emergent Dog Whistles Through Common Habitats

### v1

@@ 5.3.2 DIRECT
Table 5: Results for DIRECT Pipeline. Prec is Precision. DPR is Data Potential Recall. Best scores for across each dataset are bolded
@@
Diving into the models, LLaMa 13B does the best across Balanced and Synthetic while it fails on Realistic where LLaMa 8B does the best. This is due to Realistic being an order or two of magnitude larger than Balanced and Synthetic meaning LLaMa 13B overestimates the number of dog whistles per post in comparison to LLaMa 8B which is more conservative.
The DIRECT method outperforms all except PREDICT, which surpasses DIRECT on Balanced and Realistic datasets, while DIRECT excels on Synthetic in terms of $F_{0.5}$. The DIRECT method, though more powerful, fails in realistic settings. However, this result has caveats: DIRECT achieves significantly better DPR, identifies many dog whistles, is less conservative than PREDICT, and is not constrained by an $n$-gram threshold due to its generative nature. Overall, DIRECT achieves high DPR at the expense of precision, making it competitive but inferior to PREDICT.
@@ 6 Discussion
Conservative prediction maintains performance
The highest performing systems maintain a low number of predicted dog whistles. While it is useful for downstream tasks to have a small list of candidates, it hurts the ability of the model to find more dog whistles, creating a tricky trade off between recall and precision that fails to be satisfied by any of the current methods and EarShot.
EarShot beats all other methods
EarShot PREDICT and DIRECT both handily beat their competitors. They beat MLM/EPD by $1$-$18$ points in F-score and achieve higher precision and DPR across all scenarios. As for Word2Vec/Phrase2Vec, the gap is closer but still large. These performance highlight the possible usefulness of working towards creating a better version of EarShot.
Current model performance is not useful
Despite positive success with our method EarShot over other baselines, achieving around $1$-$25$ F-score across different datasets does not display a performance that could be used in the real world settings. More effort should be put into looking closer at designing a system that balances these, possibly leveraging the strengths of both EarShot-PREDICT and EarShot-DIRECT. In addition, it reflects on the limitations of LLMs as they might not be able to discover novel dog whistles without more help.
Future Directions
The findings of this study show the promise and limitations of current methodologies in identifying novel dog whistles. Most methods predict very few dog whistles to maintain high precision. While EarShot tries to improve on this case, its inability to balance precision and recall displays an area for future work. We propose multiple directions of improvement. First, we believe that hybrid EarShot-PREDICT and EarShot-DIRECT may prove fruitful, leveraging the strengths of both systems. Other improvements could include adding post processing to predicted dog whistles to increase the precision of methods, exploiting consensus methods through multiple LLMs, increasing the size of the models, prompting the model better methods through methods like chain of thought Wei et al. (2023) or ensembling external syntactic or semantic information. Ultimately, addressing this steep trade off between precision and recall will be vital for creating performant but applicable models.
@@ 7 Conclusion
This work lays the foundation for further inquiry into detecting new dog whistles using NLP. We develop a novel task to evaluate NLP systems in dog whistle discovery, rigorously test state of the art systems across multiple datasets, and propose a new system that surpasses the state of the art. We demonstrate that systems based on Word2Vec and BERT fail to achieve satisfactory performance on the task, whereas our EarShot system achieves respectable, though not perfect, performance across three distinct scenarios. We hope FETCH! will inspire future work on dog whistle discovery and creating difficult problems to push the boundary of what LLMs and NLP can do.
@@ Appendix A Dog Whistles Found in Each Scenario
Appendix B Prompts used in EarShot
@@
Appendix C Word2Vec/Phrase2Vec Threshold Analysis
@@
Appendix D MLM and EPD Threshold Analysis
@@
Appendix E EarShot PREDICT Threshold Analysis
E.1 EarShot BERT PREDICT Threshold Analysis
@@
E.2 EarShot LLM PREDICT Threshold Analysis

### v2

@@ 5.3.2 DIRECT
Table 5: Results for DIRECT Pipeline. Prec is Precision. DPR is Data Potential Recall. Best scores for each dataset are bolded
@@
Diving into the models, LLaMa 13B does the best across Balanced and Synthetic while it fails on Realistic where LLaMa 8B does the best as LLaMa 13B overestimates the number of dog whistles.
DIRECT outperforms all except PREDICT, which surpasses DIRECT on Balanced and Realistic datasets, while DIRECT excels on Synthetic in terms of $F_{0.5}$. DIRECT, though more powerful, fails in realistic settings. However, this result has caveats: DIRECT achieves significantly better DPR, identifies many dog whistles, is less conservative than PREDICT, and is not constrained by an $n$-gram threshold due to its generative nature.
@@ 6 Discussion
We evaluate our best EarShot systems across datasets, highlighting their strengths and limitations. We differences between models and analyze challenging dog whistles to steer potential improvements beyond EarShot. We also provide analysis of novel dog whistles found by our system in Appendix G to analyze putting our method to the test in a simulated moderation scenario.
6.1 Difference in Dog Whistles
6.1.1 Synthetic
Comparing our best model to the Unigram Word2Vec model, the Word2Vec model finds very simplistic dog whistles like “pepe”, “autogynophilia”, “13/50”, “npc”. These predictions are limited to internet slang. However, our method is able to capture more complex phraseology like “middle class”, “climate alarmists”“entitlement program” and more often used by politicians. Comparing to the bigram Phrase2Vec, the Phrase2Vec model still chooses simple internet slang like “amerimutt” and “cuckolds”. This shows that in scenarios with many dog whistles, more complex models are able to fetch more dog whistles than Word2Vec.
6.1.2 Realistic
In the Realistic setting, our method achieves high precision, distinguishing dog whistles from noise. It identifies only explicit terms like “illegal aliens,” while Word2Vec found some (e.g., “pedophilia,” “groomer”) but struggles with false positives. Moreover, Word2Vec primarily retrieves variants of seed terms (“grooming”, “pedes”), whereas our method uncovers deeper links (e.g., “mass migration” and “illegal aliens,”), highlighting stronger models in real-world scenarios, where Word2Vec fails.
6.1.3 Balanced
Every system has limitations. In the Balanced scenario, Word2Vec identifies subtler and novel dog whistles (e.g., “gibsmedats”, “we wuz kangs”, “Baltimore” “chosenites”, “africanization”, “vikangs”, “regressives”), while our method focuses on explicit slurs (“kike”, “nigger”, “faggotry.”). This likely results from our conservative threshold, which prioritizes performance, and a BERT filtering system that surfaces only the most overt dog whistles.
6.2 Difficult Dog Whistles
First, emoji-based dog whistles were difficult to find as no model detected symbols like the "OK" white power symbol, the milk emoji, or the SS logo as emoji due to the challenges in tokenizing emojis, especially in models with smaller tokenizers.
Second, context dependency and deliberate ambiguity posed significant challenges. Phrases such as “Barack Hussein Obama,” “Federal Reserve,” “New York Elite,” “Willie Horton,” and "abolish birthright citizenship" were missed due to their varied contextual usage.
Third, recency presented an issue. Newer dog whistles, like “jogger” (coined in 2020 following the shooting of Ahmaud Arbery), “save women’s sports;” (an anti-trans rallying cry), and “vegan cat” (an anti-trans dog whistle), were harder to detect due to evolving and shifting lexicons. Our best models usually defaulted to older and more known dog whistles. In the Realistic scenario, terms like “illegal immigrants” and “illegal aliens,” were detected. The Balanced scenario identified dog whistles like “ZOG,” “Baltimore,” “Sharia,” and “thug.” In the Synthetic scenario, older dog whistles such as “George Soros,” “Google,” “Rothschilds," and “White Lives Matter” were predicted. However, in the Synthetic scenario, our model did identify some newer dog whistles like “genderist” and “You Will Never Be a Woman,”, demonstrating promise in detecting emerging terms.
6.3 Future Directions
Our analysis highlights critical areas for enhancing dog whistle discovery. First, complex, nuanced, and recent phrases frequently escape discovery. Second, the dependence on hate speech classifiers tends to prioritize explicit slurs while overlooking subtler forms of harmful language. Third, the inherent precision-recall tradeoff presents substantial challenges, as evidenced by our system’s high precision and Word2Vec’s superior recall.
Additional improvements could include hybridizing Word2Vec and LLMs, post-processing noisy predictions to enhance precision, ensembling multiple LLMs, scaling model size, adding chain-of-thought reasoning Wei et al. (2023), or integrating syntactic or semantic information.
@@ 7 Conclusion
This work lays the groundwork for detecting novel dog whistles with NLP. We introduce a new evaluation task, rigorously test state-of-the-art systems across multiple datasets, and propose a superior approach. While Word2Vec and BERT-based systems under perform, our method shows promise in discovering new dog whistles but needs further development for scalability and comprehensive detection. We hope FETCH! inspires future research in dog whistle discovery and advances NLP’s capabilities.
@@ Appendix A Dog Whistles Found in Each Scenario
Appendix B Dog Whistle Examples
Dogwhistles are linguistically challenging and fairly niche use of words. We provide more examples from each dataset in context with explanations.
B.1 Balanced (Gab)
Soros
red flag warning for texas!!! gov abbott: "all americans concerned about losing our liberty to the leftist liberal agenda need to know about soros spending big in texas.
abcAnonymous Gab User
In this quote, an anonymous user on Gab informs the platform of the spending by billionaire George Soros on liberalizing reforms. Soros is a notorious boogeyman in the right-wing spaces due to his Open Society Foundations, a very influential network of non-profits that donate to liberal and progressive causes. In addition, his Jewish heritage feeds into popular antisemitic tropes about the Jews controlling the world. This dog whistle slips through many filters due to the its ability to mask possible antisemitism under the guise of criticizing billionaire donations.
Echoing
from what i read it was not primarily christian reasons jews were expelled, but for practical reasons that echo today’s reasons.
abcAnonymous Gab User
In this quote, an anonymous user on Gab informs the platform of their research on the many expulsions of the Jews from countries during history. The dog whistle here is the use of echo. While echo looks extremely benign in its usage, there are more sinister reasons to choose echo over other similar words. Echo is related to the antisemitic use of the triple parenthesis ((())) around Jewish names, topics, or names of organizations thought to be owned by Jews. These triple parentheses represent that the Jewish people’s actions "echo throughout history" hence the use of the term echo.
Gibs Me Dat
the solution to rotten tomatoes gibs me dat 100% black panthers #blaxit lawerewolf #gabfam #maga #news #auspol #speakfreely #ausfam #australia #usa
abcAnonymous Gab User
In this quote, an anonymous user on Gab insults African Americans wanting the film Black Panther to have a 100% percent approval rating on the film rating site Rotten Tomatoes. The dog whistle used here is Gibs Me Dat. Due to its extremely weird spelling and non standard usage, one might pass over it as a possible misspelling or worse dialectical usage. However, this phrase mocks African American Vernacular English, a dialect of English that African American’s often use, for being unintelligent sounding despite it having a well-formed grammar and more complex rules than Standard American English.
B.2 Synthetic (Reddit)
Baltimore
that’s what blm apparently want, no policing of the black community. they’re getting it in chicago and baltimore right now, and using their new freedom to run up the body counts to levels not seen since the drug wars of the early 90s.
abcAnonymous Reddit User
In this quote, an anonymous user on Reddit criticizes the police practices of cities like Chicago and Baltimore and the influence of the Black Lives Matter movement on these decisions. The dog whistle here is Baltimore. This dog whistle is extremely difficult to pick up because it is a regular city name. However, it is a dog whistle because it is used in a similar vein to the use of “inner cities.” Politicians reference a place that often has high African American populations like inner cities and Baltimore to disparage the people living in it. On a side note, Baltimore is a wonderful city that you should visit ☺.
Coincidences
funny how there’s been tons of rallies before but suddenly there’s all sorts of nazi imagery just so convenient for the cameras at this specific rally. probably just a coincidence i’m sure..
abcAnonymous Reddit User
In this quote, an anonymous user on Reddit believes that the appearance of Nazi imagery at a rally was not a coincidence and someone deliberately placed it there. The dog whistle here is coincidence. Coincidence often references the antisemitic trope that the Jews control the world and that many coincidences are actually Jewish conspiracies. This dog whistle slips under the radar as one could be referring to funny coincidences like being in the same vacation location as your friend at the same time or interesting coincidences like Mark Twain predicting his own death.
Family Values
holy shit… nfl is done, family values households are shutting this crap off big-time. thanks for sharing!
abcAnonymous Reddit User
In this quote, an anonymous Reddit user displays their happiness that people with family values are boycotting the National Football League due to their support of LGBTQ+ rights. The dog whistle here is family values which references the fact that the user does not believe queer life styles are normal. This dog whistle is tricky to find because family values is a catch all term that could have positive connotations but in this specific context, disguises a hatred towards queer relationships.
B.3 Realistic (Twitter)
Islamist
islamist congresswoman #ilhanomar is jeopardising american foreign policy so that she can pander to her base.
abcAnonymous Twitter User
In this quote, an anonymous Twitter user displays their hatred of Congresswoman Ilhan Omar of Minnesota for her foreign policy decisions. They deride her with the term Islamist. Islamism is an academic term which encompasses all political and religious movements that believe Islam should influence politics. However, it hides the more sinister usage of calling Ilhan Omar a terrorist, an extremist, and a traitor to the United States.
Illegals
illegals coming here by the thousands every day will drain every resource there is they will over crowd the schools
abcAnonymous Twitter User
In this quote, an anonymous Twitter user displays their distain for illegal immigrants and their supposed draining of government resources. The dog whistle here is illegals. Illegal immigrants are a diverse group of people ranging from many different places around the world. However, often when people deride illegal immigrants they often talk about Latino immigrants. This meaning often slips by filters because talking about illegal immigrants is not necessarily hate speech, but in some contexts, it is used as a cudgel against Latino people as opposed to other illegal immigrants from other countries.
Appendix C Prompts used in EarShot
@@
Appendix D Word2Vec/Phrase2Vec Threshold Analysis
@@
Appendix E MLM and EPD Threshold Analysis
@@
Appendix F EarShot PREDICT Threshold Analysis
F.1 EarShot BERT PREDICT Threshold Analysis
@@
F.2 EarShot LLM PREDICT Threshold Analysis
@@
Appendix G Novel Dog Whistles
We reviewed the top systems’ output to identify novel dog whistles in order to measure the effectiveness for dog whistle discovery. We detail 2–3 per dataset with more examples in Appendix G.
G.1 Balanced
The first dog whistle our model identified is "Wakanda," the fictional city from Marvel’s Black Panther series. It appears in contexts suggesting African Americans should return to Africa or mocking their enjoyment of the film, as in the post: “I hate Wakandans,” where it replaces the N-word. This dog whistle slips under the radar because of its benign initial contexts which could be in a discussion about the films or the comics but when used in this context it could disguise coded racism.
The second dogwhistle is “oligarchy.” Used in the post: “…the #DeepState-#Oligarchy to divide and conquer us. Don’t fall for their scam. We’re all in this together!”, it ties in with the term “deep state” which frequently implies a conspiracy about Jewish global influence. Oligarchy slips under the radar due to its multiple word senses. It could be in the post criticizing the Russian oligarchy, talking about historical oligarchies but also could hide this sense of being antisemitic.
Another novel dog whistle the model found in the Balanced scenario was brainwashing. Often, this dog whistle is used to describe either a contempt for liberal and progressive thought being supposedly being taught in public schools or Jewish control of the world. For the first meaning, the user talks about hating the supposed liberal indoctrination in schools: “true. i wouldnt let my kids go to public cesspools. and im so glad. homeschool or private school is the only answer. close the public brainwashing cesspools.” For the second meaning, explicitly the user talks about the conspiracy of Jewish global control: “nope! truth isn’t brainwashing, it’s the antidote to filthy kike lies!” (kike being a slur for a person of Jewish descent). Brainwashing could slip under the radar as it is used in describing different policies under authoritarian regimes like Nazi Germany or describing life in religious cults.
G.2 Synthetic
One novel dog whistle found from the Synthetic scenario was “zogbots”, found in the post: “…turning them into good little zogbots.” This term refers to the conspiracy theory of a Zionist Occupied Government (ZOG), suggesting Jewish control making people mindless followers.
Another novel dog whistle found in the Synthetic scenario was dindu mcnuffin. It appears in the post “..impunity while blm is mad that it now shows that dindu mcnuffin did in fact point his stolen gun at the cops/innocent bystanders/puppy.“ This dog whistle is an modification to the previously found version “dindu nuffin” which is a contraction of “did not do nothing” attempting to sound like African American Vernacular English (AAVE). This phrase mocks AAVE and references the outcry by Black communities after a Black person is a victim of police brutality where they believe that the black person did not do nothing wrong. Dindu mcnuffin could slip under the radar as its modification of dindu nuffin which could throw off classifiers or word lists for detection. In addition, its resemblance to a name also could throw off classifiers.
G.3 Realistic
Our best system did not find any new dog whistles due to its conservative threshold.